OpenStudy (anonymous):

Find the range of values P such that the straight line y=px +1 does not intersect the parabola y²=8x

OpenStudy (anonymous):

aw c'mon I've bumped twice

Miracrown (miracrown):

I have found a graphical approach... do you have use of a graphing calculator?

OpenStudy (anonymous):

I need to do it mathematically

hartnn (hartnn):

do you know how we find the intersection points ?

hartnn (hartnn):

no ?

hartnn (hartnn):

you will need to equate the 2 curves equal to each other, by squaring both sides of y= px +1 so if you equate them, you will get a quadratic in 'x'

hartnn (hartnn):

the roots of that quadratic equation are the points of intersection of those 2 curves.

hartnn (hartnn):

sin we need them to NOT intersect, we will conclude that the quadratic equation has IMAGINARY roots.

hartnn (hartnn):

and equate the discriminant part b^2-4ac as negative b^2-4ac < 0 that will give you range of p

Miracrown (miracrown):

@StudyMathlol follow Hartnn's explanation they are pretty sound!

OpenStudy (dls):

@hartnn can we find the values of P for which both of them intersect and subtract from R?

hartnn (hartnn):

try it out. if you get the same answer as that we get by my method, then YES, else no.

OpenStudy (dls):

does it sound about right anyway?

hartnn (hartnn):

yes, but why you want to do 2 steps when it can be done in 1.

OpenStudy (dls):

y=px+1 y^2=8x (px+1)^2=8x \[p^2 x^2+ 1 +2px=8x\] how can we solve this? :P

hartnn (hartnn):

you don't solve the equation you find, a,b,c and then b^2-4c <0

OpenStudy (dls):

okay got it :D

hartnn (hartnn):

@StudyMathlol did you get what i am trying to explain ?

OpenStudy (anonymous):